OpenStudy (anonymous):

Find the sixth term of a geometric sequence with t5 = 24 and t8 = 3.

OpenStudy (welshfella):

t8 = a1 r^7 = 3 t5 = a1 r^4 = 24 t8 / t5 = r^3 = 3/24 = 1/8 so r = ?

OpenStudy (welshfella):

r = cube root of 1/8

OpenStudy (anonymous):

r=1/2 right

OpenStudy (welshfella):

yes so 6th term = t5 * 1/2

OpenStudy (anonymous):

so the 6th term would be 12

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

awesome. thanks so much for your help :)

OpenStudy (welshfella):

common ratio r = 1/2

OpenStudy (welshfella):

yw